OpenStudy (anonymous):

Find the linearization L(x) of the function at a. f(x) = x^4 + 4x^2, a = −1

OpenStudy (anonymous):

I keep getting F(x) = f(x) = -12x - 8 as my asnwer but it's wrong :(

OpenStudy (anonymous):

is this right? \[L(x)=f(a)+(x-a)f'(a) \]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so we have a=-1 and\[L(x)=f(-1)+(x+1)f'(-1)\]how about this one ? make sense?

OpenStudy (anonymous):

i think u got it from here :)

OpenStudy (anonymous):

yeah but thats exactly what i do but the asnwer sumhow is wrong

OpenStudy (anonymous):

f(-1) = 5 and f'(-1) = -12. So it becomes 5 + -12(x+1) = -12x + 8

OpenStudy (anonymous):

and it keeps marking it wrong

OpenStudy (anonymous):

idk what im doing wrong here

OpenStudy (anonymous):

lemme check

OpenStudy (anonymous):

srry that shud be -12x -8

OpenStudy (anonymous):

i got ... -12x-7

OpenStudy (anonymous):

WOWW IM SUCHHH A FAILLURE...OK wow i cant beleive i got the calc right and algebra wrong

OpenStudy (anonymous):

thank u lol i really cant beleive i got this wrong

OpenStudy (anonymous):

no problem ..it happens sometime :)

OpenStudy (anonymous):

ok but this one its not about the algebra. Im just cnfused how to do it. Do u mind helping if i post it

OpenStudy (anonymous):

ah sure :)

OpenStudy (anonymous):

Thanks im posting it as a dffnt question